# Chat with AI: The Surprising Turn of Real-time Video Communication from Human to AI

### v1

@@ 4. DISCUSSIONS AND OPEN QUESTIONS
Proactive context-aware. In this paper, Artic leverages user words to achieve context awareness. It may not necessarily perform well in practice. Because it requires user words to be known before video encoding. But users may speak at any moment in the video, causing user words not to cover some segments. For example, in some benchmarks like (Lin et al., 2024; Wang et al., 2025b; Yang et al., [n. d.]), they assume that users ask questions at the end of the video. As our next step, we are building a proactive context-aware mechanism that can actively recognize important video regions even if users do not speak.
MLLM long-term memory. To minimize bitrate, Artic discards most video content irrelevant to the current chat context. This is based on the assumption that the current chat only references real-time video content. However, some MLLMs have developed long-term memory mechanisms (Wang et al., 2025a; Xiong et al., 2025; Qian et al., 2025), allowing chats to reference historical video content. Some video content, even if not relevant in the current chat context, may be needed in future chats. As our next step, we are developing a semantic layered video streaming framework. Different from SVC (Schwarz et al., 2007) that layers based on video quality, we layer by semantic correlation. The base layer contains the most important video content for the current chat context, so it must ensure low latency. The enhancement layers contain complete video details, used to offline build long-term memory, so they are not sensitive to latency.

### v2

@@ 4. DISCUSSIONS AND OPEN QUESTIONS
Proactive context-aware. In this paper, we leverage user words to achieve context awareness. It may not necessarily perform well in practice. Because it requires user words to be known before video encoding. But users may speak at any moment in the video, causing user words not to cover some segments. For example, in some benchmarks like (Lin et al., 2024; Wang et al., 2025b; Yang et al., [n. d.]), they assume that users ask questions at the end of the video. As our next step, we are building a proactive context-aware mechanism that can actively recognize important video regions even if users do not speak.
MLLM long-term memory. To minimize bitrate, the sender discards most video content irrelevant to the current chat context. This is based on the assumption that the current chat only references real-time video content. However, some MLLMs have developed long-term memory mechanisms (Wang et al., 2025a; Xiong et al., 2025; Qian et al., 2025), allowing chats to reference historical video content. Some video content, even if not relevant in the current chat context, may be needed in future chats. As our next step, we are developing a semantic layered video streaming framework. Different from SVC (Schwarz et al., 2007) that layers based on video quality, we layer by semantic correlation. The base layer contains the most important video content for the current chat context, so it must ensure low latency. The enhancement layers contain complete video details, used to offline build long-term memory, so they are not sensitive to latency.
@@
Client-side computation. Despite being optimized for mobile devices, Mobile-CLIP still incurs considerable computation. This leads to the computational resources not being fairly equalized in the comparison, because the encoders for both ours and the baseline use the default preset (medium). As future work, the baseline can adopt a more complex encoder preset (e.g., slower) to balance computational resources and achieve better video quality. Moreover, extra computational resources can also be used to explore model collaboration. For example, deploying a mobile MLLM (Jin et al., 2025; Yao et al., 2024) on the client to handle simple questions locally, while only transmitting challenging videos to the cloud-side MLLM.
Client-side tokenizer and token streaming. Is it possible to offload the video tokenizer from the server to the client and stream video tokens to the MLLM? This offers three potential gains: First, the tokenizer can serve as a powerful video compressor. For instance, MAGVIT-v2 (Yu et al., [n. d.]) achieves a better compression ratio than H.266 (Lee et al., 2023). Second, video tokens are loss-resilient. Even when 82.8% of tokens are lost, the MLLM can still maintain 98% of its original accuracy (Yao et al., 2025). On the other hand, missing tokens can be recovered at the receiver using some Masked Language Models (Li et al., 2023). Third, offloading the tokenizer can fully leverage client-side computational resources, thereby alleviating server-side pressure and increasing the number of concurrent requests. However, despite the significant potential benefits, this approach is infeasible. It is important to note that there are two types of video tokens: continuous tokens (outputs from the encoder, represented as embeddings) and discrete tokens (quantized through a codebook, represented as indices (Van Den Oord et al., 2017)). Only discrete tokens have a low bitrate (Yu et al., [n. d.]), while continuous tokens are uncompressed floating-point tensors whose bitrate is too high to stream. Discrete tokens are used only for AIGC tasks (such as text-to-video generation (Yu et al., [n. d.]; Li et al., 2025; Chen et al., 2025)), while MLLMs exclusively employ continuous tokens for video understanding (Xu et al., 2025; Li et al., 2025; Chen et al., 2025)). Although some earlier MLLMs adopted discrete tokens (Team, 2024), state-of-the-art MLLMs no longer do so due to the significant accuracy loss caused by quantization.
Circularity. The substantial gains shown in Figure 9 partially stem from circularity. Since DeViBench has ”cherry-picked” QA samples on which the MLLM makes mistakes at 200 Kbps, the baseline’s low accuracy is expected (Although different encoders were used during testing and QA selection). In fact, this is the motivation for proposing DeViBench. As discussed in §2.3, only 8% of existing QA samples exhibit errors at low bitrates and 92% can be answered correctly. This indicates that existing QA samples are too coarse-grained and lack reference to details. To bridge this gap, we need more video quality-sensitive QA samples to evaluate how quality degradation affects MLLM accuracy. In future work, we will analyze the proportion of such video quality-sensitive QA in real-world AI Video Chat applications. On the other hand, we will also test the accuracy of various MLLMs on such QA samples to demonstrate the generality of the gains.
Acknowledgements
We sincerely thank our shepherd Keith Winstein, and reviewers for their valuable feedback. This work is sponsored by the National Natural Science Foundation of China (62431017). We gratefully acknowledge the support of Key Laboratory of Intelligent Press Media Technology. Xinggong Zhang is the corresponding author (zhangxg@pku.edu.cn).